OpenStudy (anonymous):

Solve For X 5.86x+11.72=23.44 A.6 B.2 C.5.86 D.7.86

OpenStudy (igreen):

Subtract 23.44, then divide 5.86 by it..what do you get? @pootypie

OpenStudy (igreen):

Subtract 11.72*

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Which value for x makes the sentence true? 11x = 5x + 24 x = 30 x = 18 x = 4 x = 2 @iGreen

OpenStudy (anonymous):

5.86*2+11.72=23.44 yes its 2

OpenStudy (yanasidlinskiy):

Try replacing your x as one of the answer choices. For example plug in 2. 11(2) = 5(2) +24 22=10+24 Is that true? Nope! So move on to the next one!:)

OpenStudy (igreen):

\(5.86x+11.72=23.44\) Subtract 11.72 to both sides: \(5.86x = 11.72\) Divide 5.86 to both sides: \(x = 2~\Large\color{lime} \checkmark\)

OpenStudy (igreen):

11x = 5x + 24 Subtract 5x to both sides, what's 11x - 5x? @pootypie

OpenStudy (igreen):

\(11x - 5x\) is the same thing as \(11 - 5\), but you put \(x\) in your answer.

OpenStudy (anonymous):

6X

OpenStudy (igreen):

Yep, so that gives us: \(6x = 24\) \(\dfrac{\cancel6x}{\cancel 6} = \dfrac{24}{6}\) Now we divide 6 to both sides, can you divide 24 / 6?

OpenStudy (anonymous):

Yes You Get 4

OpenStudy (igreen):

Yep, that's your answer.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @pootypie

OpenStudy (anonymous):

Directions: Drag each equation to the correct location on the model. Not all equations will be used. Finish the calculation below by arranging the steps to solve for x. George bought a shirt and a pair of pants at a clothing store. The pants cost $5.71 more than the shirt, and he used a $5 off coupon to purchase the shirt. The equation below is used to determine the cost of the shirt before the coupon, s. s - $5 = $25.46 How much did the pants cost? $31.17 $36.17 $30.46 $24.75

OpenStudy (anonymous):

@iGreen